OpenStudy (anonymous):

Find the exact values of cosine, sine, and tangent for angle 225˚.

OpenStudy (anonymous):

you know the unit circle?

OpenStudy (anonymous):

no not quite

OpenStudy (anonymous):

well the cos value represents the x value and sin represents the y value on the unit circle and tan is just sin/cos. so they want you to find the x and y value at 225 degrees and y/x at 225 degrees.

OpenStudy (anonymous):

so at 225 degrees x or cos225 is \[\frac{ -\sqrt{2} }{ 2 }\]

OpenStudy (anonymous):

also sin225 which is just the y value at 225 is \[\frac{ -\sqrt{2} }{ 2 }\]

OpenStudy (anonymous):

and therefore tan225 which is just sin225/cos225 or y/x at 225 is 1

OpenStudy (anonymous):

thank you that make a whole lot more sense!

OpenStudy (anonymous):

I bet it does lol Gotta memorize that unit cicle

OpenStudy (anonymous):

i will try

OpenStudy (anonymous):

There's tricks you'll learn. good luck

OpenStudy (anonymous):

thanks you